OpenStudy (anonymous):

a bag has 5 red marbles, 2 yellow marbles, and 3 green marbles. What is the probability of drawing a red and green marble in any order (red then green or green then red) without replacing the marble you drew? How did you solve this?

OpenStudy (anonymous):

5+2+3= total marbles 5+3=red & green marbles so the number of red & green marbles/total marbles is your answer

OpenStudy (anonymous):

that would be if it was red or green, I need red and green

OpenStudy (anonymous):

8/10 is both isn't it?

OpenStudy (anonymous):

10 marbles red, then green:(5/10)(3/9) = 1/6 green, then red: (3/10)(5/9) = 1/6 combined = 1/3